Decreasing cube. Each of the three dimensions of a
cube with sides of length s centimeters is decreased by a
whole number of centimeters. The new volume in cubic
centimeters is given by
V(s)= s^3  13s^2 + 54s  72.
a) Find V(10).
b) If the new width is s  6 centimeters, then what are the
new length and height?
c) Find the volume when s=10 by multiplying the
length, width, and height.

Answer by nyc_function(2741)

V(s)= s^3  13s^2 + 54s  72.
(a) Find V(10).
To find V(10), replace s with 10 in your function and simplify. Here is the set up:
V(10) = (10)^3 - 13(10^2) + 54(10) - 72
Can you do it now?
(b) If the new width is (s  6) centimeters, then what are the
new length and height?
The volume of a cube is usually found using V = L*W*H = side^3
If the width is (s - 6), what do you think the new width could be?
(c) Find the volume when s = 10 by multiplying the
length, width, and height.
Look at part (a) of this question for your hint to part (c).
Understand?